OpenStudy (anonymous):

Please Help! An airplane travels 600 miles in the same amount of time that a car travels 200 miles. If the plane is travelling 120 mph faster than the car, find the car's speed

OpenStudy (anonymous):

t=time

OpenStudy (anonymous):

it's impossible without a time isnt it? oh god mr nalepa, the homework you give

OpenStudy (anonymous):

distance/ speed= time

OpenStudy (anonymous):

c= speed of the car \[\frac{600}{120+c}=\frac{200}{c}\]

OpenStudy (anonymous):

solve for c

OpenStudy (anonymous):

oh thank you, i thought he was trying to make an impossible question cause we werent doing our hw

OpenStudy (anonymous):

we can set them equal to each because the time it take is same for the car and the plane

OpenStudy (anonymous):

oh, now i get it, thanks a bunch

OpenStudy (anonymous):

so its equal to 200 over c or minus 200 over c?

OpenStudy (anonymous):

+200

OpenStudy (anonymous):

it is distance; can't be negative

OpenStudy (anonymous):

wait now im confused

OpenStudy (anonymous):

still confused?

OpenStudy (anonymous):

yes, how do i solve it, do i cross multiply? im not sure what that equation meant

OpenStudy (anonymous):

\[\frac{Distance}{Speed}=Time\] \[\frac{600}{120+c}=\frac{200}{c}\]

OpenStudy (anonymous):

so do i cross multiply? ho do i solve it?

OpenStudy (anonymous):

cross multiply to solve for c 600c=120*200+200c

OpenStudy (anonymous):

okay, now i understand, thank you